Find the exact value of tan (sin-1(-1/2))




1
Expert's answer
2020-08-23T17:05:41-0400

"sin^{-1}(-\frac{1}{2}) = -30^ o" Or "150^o"


"tan 150 ^o= - \frac{1}{\sqrt3}"


"tan(sin^{-1}(-\frac{1}{2})) =" "-\frac{1}{\sqrt3}"